### v2

# Luminis Stellarum et Machina: Applications of Machine Learning in Light Curve Analysis

Abstract
The rapid advancement of observational capabilities in astronomy has led to an exponential growth in the volume of light curve (LC) data, creating both opportunities and challenges for time-domain astronomy. Traditional analytical methods often struggle to fully extract the scientific value of these large and complex datasets. Machine learning (ML) algorithms are increasingly used for LC analysis, enabling classification, prediction, pattern discovery and anomaly detection. However, research in this area remains fragmented, with no comprehensive synthesis of how ML methods address the specific challenges of LC data. Key difficulties include class imbalance, noisy or sparse measurements, effective feature extraction, and limited interpretability of models. This lack of a unified overview makes it difficult for researchers to identify suitable approaches or recognize unresolved problems that require methodological advances. To address this gap, this survey systematically reviews ML techniques applied to LC analysis, outlining their principles and applications in tasks such as exoplanet detection, variable star classification and supernova identification. By clarifying the current state of the field and highlighting open challenges, this work provides guidance for future research and supports the effective integration of ML into astronomical big data studies.
Light curve classification (1954) – Variable stars (1761) – Supernovae (1668) – Neural networks (1933) – Convolutional neural networks (1938)
1 Introduction
Time domain astronomy (TDA) is a rapidly evolving field, driven by the continual discovery of new phenomena and the expansion of observational capabilities Ball & Brunner (2010); Pesenson et al. (2010). This field investigates time-varying characteristics of celestial objects using data from diverse messengers, including gravitational waves, neutrinos, and electromagnetic radiation across various photon-energy bands Vaughan (2013). With the advent of advanced wide-field, multi-epoch sky surveys such as the Sloan Digital Sky Survey (SDSS) York et al. (2000), Zwicky Transient Facility (ZTF) Graham et al. (2019), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) Kaiser et al. (2010), and the upcoming Vera C. Rubin Observatory (formerly LSST) Ivezić et al. (2019), the volume of transient discoveries has grown exponentially. For instance, ZTF generates alerts for over 100,000 events each night, while LSST is projected to surpass this by an order of magnitude. These alerts, which indicate significant flux density changes or new spatial positions, represent invaluable data streams for studying transient phenomena such as supernovae, pulsars, and gamma-ray bursts. TDA focuses on analyzing time-series data to construct LCs, which depict the variation in brightness of celestial objects over time and are fundamental to understanding the physical processes driving these phenomena.
However, harnessing these vast data streams requires addressing significant challenges. It is important for astronomers to take into account the possible uncertainties and biases that may arise from observational methods as well as from data analysis procedures. Observational uncertainties stem from limitations in instrumentation and data collection. Analytical uncertainties, on the other hand, emerge from the complexities of modeling physical processes and the numerical methods used to approximate them. Mitigating biases requires developing comprehensive models that incorporate diagnostics from multiple perspectives. This necessitates detailed multi-physics simulations that combine data from various sources and rely on cutting-edge computational capabilities. As a result, TDA not only facilitates the study of astrophysical phenomena but also serves as a bridge to advancing fundamental physics, making it a cornerstone of modern astronomical research.
ML has become an essential tool for automating data classification, archiving, and retrieval Fluke & Jacobs (2020); Baron (2019). Unlike traditional feature extraction methods, which require domain expertise, deep learning (DL) automatically learns data representations. As a subset of ML, DL has gained widespread popularity and has been successfully applied in various domains, including speech recognition, image processing, and natural language processing Nassif et al. (2019); Deng & Li (2013); Rodellar et al. (2018); Mishra & Celebi (2016); Ussipov et al. (2024a, b); Otter et al. (2020). In astronomy, ML has emerged as a powerful tool for automating the classification and analysis of LCs, enabling researchers to extract insights from vast datasets.
Despite its growing adoption, several persistent challenges limit the effectiveness of ML for LC analysis. Class imbalance occurs because some astronomical sources, such as rare transients, are vastly underrepresented compared to more common classes, which can bias models toward majority classes and reduce sensitivity to rare events. Noise arises from observational uncertainties, instrumental limitations, or environmental effects, often degrading model performance and leading to spurious detections. Interpretability is another major challenge, as many ML and DL methods function as “black boxes,” making it difficult to relate predictions to physical processes or diagnose errors. In addition, biases and uncertainties can stem from selection effects in survey design, incomplete or mislabeled training data, and systematic measurement errors; these issues can propagate through the ML pipeline and affect generalizability. Addressing these challenges is crucial for building models that not only produce accurate predictions but also offer trustworthy scientific insights.
This article provides a comprehensive overview of ML algorithms in LC analysis, covering applications such as exoplanet detection, variable star classification, and supernova identification. By presenting a structured examination of ML techniques in this domain, the review aims to foster a deeper understanding of current methodologies and identify chanllenges that should be adressed in the near feature.
The rest of the paper is organized as follows. The methodology is presented in Section 2. Section 3 provides an overview of major photometric data surveys relevant to LC analysis. Section 4 introduces fundamental ML concepts and architectures. Section 5 explores key applications of ML techniques to LC analysis. Section 6 outlines current challenges and open research questions in the field. Finally, Section 7 concludes with a summary and outlook.
2 Review Methodology
To compile this review, we followed the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines Moher et al. (2009). The overall review methodology is illustrated in Figure 1. A comprehensive search was conducted across major scientific databases, including IEEE Xplore, SpringerLink, Elsevier, Web of Science, and Google Scholar, using Boolean combinations of keywords such as “light curve classification”, “light curve machine learning”, “variable star classification”, “supernova classification”, and “exoplanet detection”. The initial search identified 2,780 records, of which 680 duplicates were removed. Title and abstract screening excluded an additional 1,951 records, followed by the removal of 16 records after full-text eligibility assessment. Ultimately, 133 studies were included in this review.
The review covers publications from 2005 to 2025. Studies were included if they (1) were peer-reviewed, (2) applied ML methods to real or simulated LC data, and (3) were relevant to tasks such as variable star classification, exoplanet detection, or transient identification. Non-English papers, duplicate entries, and studies lacking experimental validation or methodological clarity were excluded.
Figure 1: PRISMA flowchart of review methodology.
3 Photometric Survey Datasets
Over the past few decades, large-scale photometric surveys have transformed time-domain astronomy, providing vast, multi-wavelength datasets that are readily accessible through online databases. These surveys have been essential in detecting and classifying variable stars, discovering exoplanets, and identifying transient astronomical events. As shown in Figure 2, the data volume from astronomical surveys has grown exponentially, approximately doubling every 16 months. The combination of different survey datasets enables comprehensive studies of stellar variability across various timescales and wavelengths, driving both traditional astrophysical research and modern ML applications. This rapid data growth has created a strong demand for automated and scalable analysis methods, especially those based on ML.
Figure 2: The total data volume produced by various astronomical surveys throughout their lifetimes. The data reveals that the volume of astronomical survey data doubles approximately every 16 months. The data and figure are sourced from Kremer et al. (2017); Smith & Geach (2023).
Understanding the properties of these surveys is essential for ML applications, since survey-specific factors such as cadence, photometric precision, time coverage, and instrumental effects directly affect the structure and quality of LC data. In addition, biases and uncertainties introduced during survey design or data processing can influence ML model performance, generalization, and reliability. Summarizing major photometric surveys provides important context for selecting appropriate ML techniques and for understanding the practical limitations of real-world astronomical data.
Photometric surveys can be broadly categorized into ground-based and space-based missions. Ground-based surveys provide wide-field and high-cadence monitoring, often covering extensive regions of the sky over long time periods. In contrast, space-based missions offer high-precision, uninterrupted observations unaffected by atmospheric distortions, enabling the detection of minute variations in stellar brightness. Table LABEL:table1 provides an overview of major ground-based and space-based photometric surveys, including their active years, observed bands, sky coverage, and references.
Table 1: Overview of major photometric surveys, including their active years, observed bands, and sky coverage.
Surveys Active years Bands Area Reference
Hipparcos / Tycho 1989-1993 $B_{T}$, $V_{T}$ MW Perryman et al. (1997); Høg et al. (1997)
MAssive Compact Halo Objects (MACHO) 1992-1999 V, R MW, LMC, SMC Alcock et al. (2000)
Optical Gravitational Lensing Experiment (OGLE) 1992-present V, I MW, LMC, SMC Udalski et al. (2015)
All Sky Automated Survey (ASAS) 1997-present V, I All sky Pojmanski (1997)
Two Micron All-Sky Survey (2MASS) 1997-2001 J, H, $K_{s}$ All sky Skrutskie et al. (2006)
Lincoln Near-Earth Asteroid Research (LINEAR) 1998-2015 Unfiltered Wide field Stokes et al. (2000)
Robotic Optical Transient Search Experiment (ROTSE) 1998-present Unfiltered All sky Akerlof et al. (2003)
Sloan Digital Sky Survey (SDSS) 1998-present u, g, r, i, z Stripe 82 York et al. (2000)
Northern Sky Variability Survey (NSVS) 1999-2004 Unfiltered Northern sky Woźniak et al. (2004)
Catalina Real-Time Survey (CRTS) 2003-present V All sky Drake et al. (2009)
Hungarian Automated Telescope Network (HATnet) 2003-present r Wide field Bakos et al. (2004)
Wide Angle Search for Planets (WASP / SuperWASP) 2004-present Optical All sky Pollacco et al. (2006)
UKIRT Infrared Deep Sky Survey (UKIDSS) 2005-2014 Z, Y, J, H, K Wide field Lawrence et al. (2007)
Convection, Rotation and Planetary Transits (CoRoT) 2006-2013 Unfiltered MW Barge et al. (2008)
Kepler mission 2009-2018 Unfiltered MW Koch et al. (2010)
Wide field Infrared Survey Explorer (WISE) 2009-present $W_{1}$, $W_{2}$, $W_{3}$, $W_{4}$ Wide field Wright et al. (2010)
VISTA Variables in the Vía Láctea (VVV) 2010-2016 Z, Y, J, H, $K_{s}$ Wide field Minniti et al. (2010)
High Cadence Transit Survey (HiTS) 2013-2015 u, g, r, i Wide field Förster et al. (2016)
Gaia 2013-present BP, RP All sky Prusti et al. (2016)
All Sky Automated Survey for SuperNovae (ASAS-SN) 2014-present V, g All sky Kochanek et al. (2017)
Next Generation Transit Survey (NGTS) 2015-present I Wide field Wheatley et al. (2018)
Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) 2010-present g, r, i, z, y Wide field Magnier et al. (2013)
Transiting Exoplanet Survey Satellite (TESS) 2018-present TESS-band All sky Ricker et al. (2015)
Zwicky Transient Facility (ZTF) 2018-present g, r, i Northern sky Bellm et al. (2018)
James Webb Space Telescope (JWST) 2021-present IR Targeted fields Gardner et al. (2023)
Wide Field Survey Telescope (WFST) 2023-present u, g, r, i, z, w Northern sky Lou et al. (2016)
Vera C. Rubin Observatory (formerly LSST) 2025 (expected) u, g, r, i, z, y Wide field Ivezić et al. (2019)
PLAnetary Transits and Oscillations of Stars (PLATO) 2026 (expected) B, V, R, I Wide field Rauer et al. (2014)
Ultraviolet Transient Astronomy Satellite (ULTRASAT) 2027 (expected) NUV All sky Shvartzvald et al. (2024)
Roman Space Telescope (formerly WFIRST) 2027 (expected) IR Wide field Spergel et al. (2015)
4 Machine learning fundamentals
To demonstrate the applications of ML in LC analysis, this section provides a concise overview of core ML concepts.
Figure 3: Overview of main ML approaches.
Figure 4: Basic process of ML. Cylinder represents database, rectangles indicate data processing steps, and the rectangle with a folded corner denotes knowledge extraction.
ML enables systems to autonomously learn patterns from data and improve decision-making through experience, mirroring aspects of human cognition. Unlike traditional astronomical programming, which relies on explicit physical rules, ML algorithms derive implicit relationships directly from observational data. This data-driven approach offers flexibility for solving complex, nonlinear problems that defy conventional analytical methods Rodr´ıguez et al. (2022); Kembhavi & Pattnaik (2022); Sen et al. (2022).
ML algorithms can be broadly categorized into supervised and unsupervised methods, often referred to as predictive and descriptive, respectively. These approaches may also be combined to form semi-supervised methods. Figure 3 provides a visual taxonomy of these approaches. Supervised algorithms learn mappings between input features and predefined target variables using labeled training data curated by domain experts Zucker & Giryes (2018); Delli Veneri et al. (2019); Krone-Martins et al. (2018); Mahabal et al. (2019); D’Isanto et al. (2018); Norris et al. (2019). Unsupervised methods, conversely, autonomously identify hidden structures or relationships within unlabeled datasets. These techniques are typically divided into three subcategories: clustering (grouping similar data points), dimensionality reduction (extracting salient features), and anomaly detection (identifying outliers) Hocking et al. (2015); Nun et al. (2016); Gianniotis et al. (2016); Baron & Poznanski (2017); Reis et al. (2018a, b). Anomaly detection holds particular promise for astronomical research, as it enables discovery of rare or unexpected phenomena within existing observational datasets.
Another major category is reinforcement learning, which is based on learning through interaction with an environment to maximize cumulative rewards. However, it is excluded from the scope of this review, as its application to LC analysis remains limited due to the lack of interactive, feedback-driven settings in typical observational astronomy workflows.
As an interdisciplinary field, ML integrates principles from statistics, optimization theory, and information science. Its primary focus lies in developing adaptive systems that simulate human learning processes to iteratively acquire knowledge, refine skills, and optimize performance. Figure 4 illustrates the iterative workflow of a typical ML system. Knowledge extraction from data can be effectively achieved using established methodologies such as the Knowledge Discovery in Databases (KDD) process Fayyad et al. (1996) and the Cross-Industry Standard Process for Data Mining (CRISP-DM) Wirth & Hipp (2000).
4.1 Supervised Learning
Supervised learning (SL) operates under guidance, where labeled data provides the necessary supervision for model training. In this framework, class labels or continuous target values act as references, enabling the model to learn mappings between inputs and outputs. The basic flow of SL is shown in Figure 5. SL problems are typically divided into classification and regression. Classification involves predicting discrete class labels (e.g., variable stars based on their LC), while regression involves predicting continuous values (e.g., estimating a star’s period or temperature from its LC). Both approaches have been applied to LC analysis: classification is widely used for star type identification, supernova classification and exoplanet detection Akhmetali et al. (2024); Mahabal et al. (2017); Valizadegan et al. (2022); Pearson et al. (2018); Mitra et al. (2023, 2024); Abylkairov et al. (2025); Bissekenov et al. (2024), whereas regression is used for parameter estimation and LC detrending Hinners et al. (2018); Thieler et al. (2016); Aigrain et al. (2016); Paunzen & Vanmunster (2016); Carter & Winn (2009).
Figure 5: Basic process of SL. Cylinder represents database, rectangles indicate data processing steps, parallelograms denote datasets, and the rounded rectangle shows the final output.
These algorithms optimize their performance by minimizing a cost function, which quantifies the discrepancy between predicted and actual values. The greater the deviation, the more challenging it becomes to achieve accurate predictions. Effective learning depends on a well-curated dataset with precise class labels, as a larger and higher-quality training set facilitates smoother optimization and improves model accuracy.
Figure 6: Overview of SL algorithms. (a) RF: An ensemble of DTs with majority voting or averaging. (b) MLP: A feedforward neural network with input, hidden, and output layers. (c) RNN: Processes sequential data using recurrent connections (denoted as red lines) between hidden layers. (d) LSTM: A specialized RNN with memory cells to capture long-term dependencies, utilizing recurrent connections (denoted as red lines) for information flow. (e) SVM: Identifies an optimal hyperplane with maximum margin for class separation. (f) KNN: Classifies data based on the majority vote of the $k$ nearest neighbors. (g) CNN: Extracts hierarchical features using convolutional and pooling layers, followed by fully connected layers for classification.
Decision Trees (DTs) are hierarchical, non-parametric models that recursively partition the feature space by applying conditions on features at internal nodes Quinlan (1986). The model is defined by this tree structure, with nodes, branches, and leaves representing features, decisions, and outcomes, respectively. The learning algorithm trains the tree by selecting splits that minimize impurity measures such as entropy for classification or mean squared error (MSE) for regression:
$\text{Entropy}(R)=-\sum_{k=1}^{K}p_{k}\log_{2}p_{k},$
$\text{MSE}(R)=\frac{1}{|R|}\sum_{i\in R}(y_{i}-\bar{y}_{R})^{2},$
where $p_{k}$ is the probability of class $k$ in region $R$, and $\bar{y}_{R}$ is the mean of the target values in $R$. The algorithm also includes strategies to mitigate overfitting and bias, such as pruning or limiting tree depth, and handles class imbalance using cost-sensitive splitting or balanced sampling. DTs naturally quantify uncertainty via the class probability distribution at leaf nodes. DTs are used for star–galaxy separation Vasconcellos et al. (2011), detecting active objects Zhao & Zhang (2008), redshift estimation Suchkov et al. (2005), and classifying stars, galaxies, and AGNs Golob et al. (2021).
Random Forests (RFs) are ensembles of $B$ DTs, which together define the model Breiman (2001). Each tree partitions the feature space into regions corresponding to its leaf nodes, where a region $R$ contains all samples reaching that leaf. The learning algorithm trains each tree on bootstrap samples with random feature subsets. Predictions are then aggregated from all trees via majority voting for classification or averaging for regression:
$\hat{y}=\text{majority vote}\left(\{f_{b}(x)\}_{b=1}^{B}\right)\quad\text{or}\quad\hat{y}=\frac{1}{B}\sum_{b=1}^{B}f_{b}(x),$
Randomization reduces correlation among trees and enhances robustness. Uncertainty can be estimated using out-of-bag (OOB) predictions as an internal validation measure. Random feature selection mitigates bias and overfitting, while class imbalance is handled through balanced class weights. Figure 6(a) illustrates a typical RF architecture. RFs are widely used for supervised tasks such as variable star or supernova classification Jenkins et al. (2012); Kim & Bailer-Jones (2016); Gomez et al. (2020), and can also be applied in unsupervised settings to estimate object similarity Reis et al. (2018a, b).
Artificial Neural Networks (ANNs) are nonlinear, adaptive computational models composed of interconnected units called neurons Haykin (1994). The model is defined by the network architecture, specifying layers, neurons, and connectivity. Each neuron computes an output:
$y=\phi(w^{T}x+b),$
where $\phi$ is an activation function, $w$ is a vector of connection weights, $x$ is the input vector, and $b$ is a bias term. The learning algorithm, such as backpropagation, adjusts the weights to minimize prediction error. Inspired by biological neural networks, ANNs simulate learning and decision-making by adapting connections based on input data, making them highly effective for complex pattern recognition and predictive modeling in high-dimensional domains. Class imbalance is addressed using weighted loss functions, while uncertainty can be approximated via dropout or Bayesian layers. ANNs are used for classification of galaxy spectra Folkes et al. (1996), modeling Spectral Energy Distribution (SED) of galaxies Silva et al. (2011), probing the opacity of Universe to cosmic gamma rays Singh et al. (2022), and prediction of the atmospheric fundamental parameters from stellar spectra Azzam et al. (2021).
Multilayer Perceptrons (MLPs) are feedforward ANNs composed of an input layer, one or more fully connected hidden layers, and an output layer. The model is defined by the architecture specifying the number of layers, neurons per layer, and activation functions. Each hidden layer computes:
$\mathbf{h}_{l}=\sigma(\mathbf{W}_{l}\mathbf{h}_{l-1}+\mathbf{b}_{l}),$
where $\sigma$ is an activation function, typically the rectified linear unit (ReLU) Pal & Mitra (1992), $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ are the weight matrix and bias vector for layer $l$, and $\mathbf{h}_{0}=\mathbf{x}$ is the input vector. The learning algorithm, such as backpropagation, adjusts the weights to minimize loss functions such as cross-entropy (classification) or MSE (regression). Uncertainty can be modeled using dropout or Bayesian layers. Regularization techniques such as L2 weight decay reduce bias and prevent overfitting. Class imbalance is typically addressed by applying class-weighted loss functions or oversampling minority classes. Figure 6(b) illustrates a typical MLP architecture. MLPs are widely used in astronomy for tasks such as star–galaxy classification and distinguishing galaxies, stars, and quasars Andreon et al. (2000); Abd-elaziem & Soliman (2023); Zeraatgari et al. (2024).
Recurrent Neural Networks (RNNs) are neural networks with recurrent connections, defined by an architecture that specifies the number of layers, hidden units, and the recurrence structure for sequential data Bengio et al. (1994). The hidden state $h_{t}$ at time step $t$ is computed as:
$h_{t}=\sigma(W_{h}h_{t-1}+W_{x}x_{t}+b),$
where $x_{t}$ is the input at time $t$, $W_{h}$ and $W_{x}$ are weight matrices, $b$ is a bias vector, and $\sigma$ is an activation function (often $\tanh$ or ReLU). The learning algorithm, such as backpropagation through time (BPTT), adjusts the weights to capture temporal dependencies for sequence modeling and prediction Sutskever et al. (2011); Mikolov et al. (2014). Uncertainty is often estimated using techniques like dropout. Bias and instability due to vanishing or exploding gradients are mitigated by gradient clipping and batch normalization. Class imbalance is addressed through sequence-level weighting or synthetic time-series generation. Figure 6(c) illustrates a typical RNN architecture. RNNs are widely used in astronomy for time-series tasks, including supernova and variable star classification Charnock & Moss (2017); Naul et al. (2018), gravitational wave denoising Shen et al. (2019), periodic variable detection Tsang & Schultz (2019), and real-time transient discovery Möller & de Boissière (2020); Muthukrishna et al. (2019).
Convolutional Neural Networks (CNNs) are DL models designed for grid-like data such as images and time series Li et al. (2021); O’shea & Nash (2015); Yamashita et al. (2018). A CNN model consists of learnable convolutional filters $W$ applied to local regions $X$ of the input, extracting spatially hierarchical features:
$(W*X)_{i,j}=\sum_{m,n}W_{m,n}X_{i+m,j+n},$
Pooling layers reduce spatial dimensions, improving robustness to translations and distortions, while fully connected layers aggregate extracted features for final predictions. The corresponding learning algorithm adjusts the filter weights and biases via backpropagation and gradient descent. Uncertainty can be estimated using dropout or test-time augmentation (TTA). In TTA, multiple augmented versions of the input (e.g., rotations, flips, or shifts) are passed through the trained CNN at inference time, and the predictions are aggregated, for example via averaging or voting. The variability of predictions across these augmentations provides an empirical measure of uncertainty, while the aggregation improves robustness by reducing sensitivity to input transformations and noise. Bias is further reduced through data augmentation during training, which improves model invariance. Class imbalance is addressed using loss functions such as focal loss, which downweight well-classified examples. Figure 6(g) illustrates a typical CNN architecture. CNNs excel in astronomy tasks such as galaxy morphology classification Dieleman et al. (2015); Aniyan & Thorat (2017), variable star classification Hon et al. (2017), supernova identification Brunel et al. (2019); Pasquet et al. (2019), photometric redshift estimation Hoyle (2016), cosmological parameter inference Ntampaka et al. (2020), and strong gravitational lensing detection Lanusse et al. (2018).
Long Short-Term Memory (LSTM) networks are a specialized type of RNN, defined by an architecture that includes memory cells and gating mechanisms (input, forget, and output gates) to overcome vanishing and exploding gradient problems Hochreiter & Schmidhuber (1997); Le et al. (2015); Gers et al. (2000). The cell state $c_{t}$ is updated as:
$c_{t}=f_{t}\odot c_{t-1}+i_{t}\odot\tanh(W_{c}[h_{t-1},x_{t}]+b_{c}),$
where $\odot$ denotes element-wise multiplication. The learning algorithm, typically BPTT, adjusts the weights of the gates and hidden units to capture long-term dependencies in sequential data. Figure 6(d) illustrates a typical LSTM architecture. LSTMs are used for anomaly detection in LCs Zhang & Zou (2018), classification of solar radio spectra Xu et al. (2019), detection of core-collapse supernovae Iess et al. (2023), and solar flare prediction Sun et al. (2022).
Support Vector Machines (SVMs) are SL models that define a hyperplane separating classes in an $N$-dimensional space. The model is characterized by the weight vector $w$, bias $b$, and, for non-separable data, slack variables $\xi_{i}$. The learning algorithm then identifies the optimal hyperplane by solving the optimization problem:
$\min_{w,b,\xi}\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{N}\xi_{i},$
$\text{subject to}\quad y_{i}(w^{T}x_{i}+b)\geq 1-\xi_{i},\quad\xi_{i}\geq 0,\quad i=1,\dots,N,$
where $C$ is a regularization parameter controlling the trade-off between margin size and classification error. Nonlinear decision boundaries are handled via kernel functions such as the radial basis function (RBF) kernel. Uncertainty can be estimated using Platt scaling Platt et al. (1999), and class imbalance is addressed by adjusting class weights or using synthetic oversampling methods. Figure 6(e) illustrates a typical SVM separating hyperplane. SVMs are widely applied in astronomy for star-galaxy separation Kovács & Szapudi (2015), identification of galaxies Krakowski et al. (2016), classification of strong gravitational lenses Hartley et al. (2017), identification of pre-main-sequence stars Ksoll et al. (2018), and variable stars Pashchenko et al. (2018).
$k$-Nearest Neighbors (KNN) are non-parametric, instance-based models that define a prediction as a function of the $k$ closest training samples according to a distance metric, typically Euclidean distance Mucherino et al. (2009); Keller et al. (1985). For a query point $x$, the model output is given by:
$\hat{y}(x)$ $=\text{majority vote}\left(\{y_{i}\mid x_{i}\in\mathcal{N}_{k}(x)\}\right),$
$\mathcal{N}_{k}(x)$ $=\left\{\text{the }k\text{ nearest neighbors of }x\right\},$
The learning algorithm then determines the neighbors at prediction time using the chosen distance metric. As lazy learners, KNNs require no explicit training but can be computationally expensive for large datasets. Uncertainty in KNN predictions can be quantified using distance-weighted voting, where closer neighbors have higher influence on the predicted class, or by computing confidence scores based on the relative distances of neighbors from the query point. Feature scaling (normalization or standardization) is important to reduce bias in distance calculations. Class imbalance is addressed by adapting $k$ or weighting votes to emphasize minority classes. Figure 6(f) illustrates a typical KNN structure. KNN is widely applied in astronomy for classifying celestial objects Li et al. (2008), photometric redshift estimation Zhang et al. (2013), star–galaxy separation Machado et al. (2016a), and solar flare prediction Li et al. (2007).
4.2 Unsupervised Learning
Unsupervised learning (UL) is a ML approach that trains on unlabeled data, identifying underlying patterns and structures without predefined labels. Unlike supervised methods, which rely on labeled examples, unsupervised algorithms autonomously detect relationships, group similar data points, and uncover anomalies without external guidance. The basic flow of UL is shown in Figure 7.
UL encompasses a broad range of statistical techniques for data exploration, including clustering, dimensionality reduction, visualization, and anomaly detection. These methods are especially valuable in scientific research, as they enable the discovery of hidden patterns and the extraction of new insights from large datasets.
Figure 7: Basic process of UL. Cylinder represents database, rectangles indicate data processing steps and rounded rectangle shows the final output.
Figure 8: Overview of UL algorithms. (a) K-means: Partitions data into $k$ clusters by minimizing the variance within each cluster. (b) HC: Hierarchical clustering builds a hierarchy of clusters. (c) PCA: Reduces dimensionality by projecting data onto orthogonal axes that maximize variance, with the first principal component shown. (d) t-SNE: Embeds high-dimensional data into a low-dimensional space while preserving local structures. (e) AE: An autoencoder consisting of an encoder and decoder, used to learn a low-dimensional latent representation of the input data. (f) SOM: A self-organizing map that projects high-dimensional data onto a low-dimensional grid while preserving topological properties.
K-means is a widely used clustering model valued for its simplicity and efficiency. The model partitions data into $K$ clusters, each represented by a centroid $\mu_{k}$, and aims to minimize the within-cluster variance:
$\min_{\{C_{k}\}}\sum_{k=1}^{K}\sum_{x_{i}\in C_{k}}\|x_{i}-\mu_{k}\|^{2},$
The corresponding K-means learning algorithm initializes $K$ cluster centroids randomly, then iteratively assigns points to the nearest centroid and updates centroids until convergence MacQueen (1967); Hartigan & Wong (1979); Sinaga & Yang (2020). Uncertainty is addressed through soft clustering variants, while bias is mitigated via careful centroid initialization. K-means is sensitive to class imbalance but can be adapted using density-based weighting. Figure 8(a) illustrates how K-means works, showing the iterative process of centroid updates and cluster assignments. K-means is used for classifying stellar spectra Almeida & Prieto (2013), distinguishing stellar clusters based on chemical abundances Garcia-Dias et al. (2019), determining membership probabilities in open clusters El Aziz et al. (2016), and classifying eclipsing binary LCs Modak et al. (2020).
Hierarchical clustering (HC) is a popular clustering model that builds a hierarchy of clusters Ward Jr (1963). There are two main types of HC: Agglomerative Hierarchical Clustering (AHC) and Divisive Hierarchical Clustering (DHC). AHC, also known as the ”bottom-up” approach, models each data point as an individual cluster and defines a merging criterion based on a chosen distance metric. DHC, or the ”top-down” approach, models all data points initially as a single cluster and defines a splitting criterion to recursively divide clusters until each point becomes its own cluster Murtagh & Contreras (2012); Johnson (1967). The HC learning algorithm builds the cluster hierarchy using linkage criteria:
$d(C_{i},C_{j})=\begin{cases}\min_{x\in C_{i},\,y\in C_{j}}d(x,y)&\text{(single linkage)},\\ \max_{x\in C_{i},\,y\in C_{j}}d(x,y)&\text{(complete linkage)},\end{cases}$
Cophenetic correlation quantifies uncertainty, and linkage choice affects bias. Figure 8(b) illustrates AHC, showing the process of merging clusters in a ”bottom-up” manner. Applications of HC include identifying galaxy cluster substructures Yu et al. (2015), detecting superclusters of galaxies Santiago-Bautista et al. (2020), exploring the hierarchical structure of open clusters Yu et al. (2020), and clustering astronomical spectra Tian et al. (2025).
Principal Component Analysis (PCA) is a dimensionality reduction model that transforms high-dimensional data into a lower-dimensional representation while retaining as much relevant information as possible Wold et al. (1987). The model identifies orthogonal directions, called principal components, which capture the maximum variance in the dataset. Projecting data onto these components reduces the number of features, mitigates the curse of dimensionality, enhances computational efficiency, and improves interpretability Abdi & Williams (2010); Jolliffe & Cadima (2016).
The PCA learning algorithm computes the eigenvectors of the covariance matrix and projects data onto the top-$k$ eigenvectors:
$X^{\prime}=XV_{k},\quad\Sigma=X^{T}X,$
where $V_{k}$ contains the top-$k$ eigenvectors. Reconstruction error estimates uncertainty, and standardization prevents feature scale bias. Figure 8(c) demonstrates the PCA process. PCA is applied in astronomy for searching galaxy-scale strong lenses Joseph et al. (2014), exploring abundances in different environments Ting et al. (2012), and analyzing LCs of variable stars Deb & Singh (2009).
t-Distributed Stochastic Neighbor Embedding (t-SNE) is a dimensionality reduction model designed to represent high-dimensional data in two or three dimensions while preserving local similarity relationships Van der Maaten & Hinton (2008). The model maps similar objects close together and dissimilar objects farther apart, capturing non-linear relationships and revealing clusters and intricate patterns in complex datasets Van Der Maaten (2014); Kobak & Berens (2019). The t-SNE learning algorithm minimizes the Kullback-Leibler divergence between probability distributions in the high- and low-dimensional spaces:
$KL(P\|Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)},$
where the perplexity parameter balances local and global structure, affecting uncertainty. Careful tuning is required to avoid initialization bias. Figure 8(d) illustrates the t-SNE process. t-SNE is applied in astronomy to identify SED outliers Quispe-Huaynasi et al. (2025), reduce spectral information Traven et al. (2017), quantify similarity between planetary systems Alibert (2019), and separate Type-2 AGN from HII galaxies Zhang et al. (2020).
Autoencoder (AE) is an ANN model designed to learn efficient low-dimensional representations of input data, commonly used for tasks such as compression, dimensionality reduction, and visualization Yang & Li (2015); Gianniotis et al. (2015, 2016). The model consists of two main components: an encoder, which compresses the input data into a lower-dimensional latent representation, and a decoder, which reconstructs the original data from this compressed form. The AE learning algorithm optimizes network weights by minimizing the reconstruction error, typically measured as the squared difference between the input and reconstructed output:
$\mathcal{L}=\|x-g_{\theta}(f_{\phi}(x))\|^{2},$
where $f_{\phi}$ encodes and $g_{\theta}$ decodes. Dropout layers estimate uncertainty, while regularization prevents overfitting. Figure 8(e) shows the structure of an AE. AEs are applied in astronomy to estimate gravitational wave parameters Gabbard et al. (2022), identify strong gravitational lenses Cheng et al. (2020), estimate atmospheric parameters Yang & Li (2015), and detect anomalies in cosmic ray variations Mandrikova & Mandrikova (2022).
Self-organizing map (SOM), also known as a Kohonen map Kohonen (1982), is an unsupervised ANN model that creates a low-dimensional, typically two-dimensional, representation of high-dimensional input data. The model represents the output neurons with weight vectors that act as prototypes or templates, capturing the essential features of the input dataset. This structure allows SOMs to effectively map and visualize high-dimensional data in a more interpretable, lower-dimensional form while preserving the input topology. The SOM learning algorithm organizes the map by minimizing the quantization error:
$Q=\sum_{i}\|x_{i}-w_{c(i)}\|^{2}.$
where $w_{c(i)}$ is the winning neuron. Quantization error measures uncertainty, while neighborhood preservation reduces topological bias. Figure 8(f) illustrates the SOM process. SOMs are applied in astronomy to estimate galaxy photometric redshift probability density functions (P
4.3 Semi-supervised Learning
Semi-supervised learning (SSL) is a ML approach that integrates elements of both SL and UL. It leverages a small amount of labeled data alongside a large pool of unlabeled data, making it particularly useful when labeled data is scarce but unlabeled data is abundant Chapelle et al. (2009).
The SSL process begins with dataset collection, comprising both labeled and unlabeled data, followed by cleaning and preprocessing to ensure consistency. The model is initially trained on the labeled data to establish a foundational understanding of the task. It then refines its performance by incorporating information from the unlabeled data, improving overall accuracy. The general workflow of SSL is illustrated in Figure 9.
SSL can be categorized into inductive and transductive learning methods Van Engelen & Hoos (2020). Inductive learning aims to develop a generalized model capable of making predictions on unseen data. It utilizes both labeled and unlabeled data during training to enhance generalization. In contrast, transductive learning focuses on predicting labels solely for the specific unlabeled data available during training, without aiming for broader generalization. Transductive methods often exploit the inherent structure of the unlabeled data, such as relationships between data points, to improve prediction accuracy.
SSL offers several advantages, including more efficient data utilization by leveraging both labeled and unlabeled data, reducing the cost associated with manual labeling. Additionally, it enhances model performance by capturing structural patterns in unlabeled data. By bridging the gap between SL, which relies on labeled data, and UL, which identifies patterns without predefined labels, SSL provides a robust framework for learning in scenarios with limited labeled datasets.
[…]
Figure 9: Basic process of SSL. Cylinders represent databases, rectangles indicate data processing steps, and the rounded rectangle shows the final output.
4.4 Uncertainty and Bias
In ML models, uncertainty arises from two primary sources. Epistemic uncertainty (model uncertainty) reflects the lack of knowledge due to limited or unrepresentative training data and can be reduced by collecting additional observations or using ensemble or Bayesian methods. For example, a CNN trained on a limited set of supernova LCs may be uncertain when classifying rare or novel transients. Aleatoric uncertainty (data uncertainty) originates from inherent noise in measurements, such as photometric errors or instrumental noise, and cannot be reduced by more data. Bias can also occur systematically due to model assumptions or dataset composition; for instance, overrepresentation of bright galaxies may lead to misclassification of faint objects. Characterizing both uncertainty and bias is crucial in astronomy, as it informs the reliability of classifications, the confidence in parameter estimates, and the prioritization of follow-up observations. Techniques such as Monte Carlo dropout, Bayesian neural networks, and careful data augmentation are employed to quantify and mitigate these effects across the models.
4.5 Evaluation Metrics and Validation
4.5.1 Evaluation Metrics
Evaluation metrics are essential tools in ML that quantitatively assess a model’s performance. They play a crucial role in optimizing hyper-parameters, evaluating model effectiveness, selecting the most relevant features, and comparing different ML algorithms. These metrics are computed during both the validation and testing phases, where the trained model is applied to previously unseen data, and its predictions are compared against actual target values to measure accuracy and reliability.
In regression problems, where the target variable is continuous, model performance is commonly evaluated using metrics such as the Mean Absolute Error (MAE) and the Mean Squared Error (MSE). The MAE is defined as:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$
where $n$ is the number of instances in the validation or test set, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the model’s prediction. The MSE is given by:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}.$
Moreover, additional evaluation metrics are also found in the literature, such as the Normalized Median Absolute Deviation, the Continuous Ranked Probability Score, and the Probability Integral Transform (see e.g., D’Isanto & Polsterer (2018); D’Isanto et al. (2018)).
In classification tasks, where the target variable is discrete, commonly used evaluation metrics include Classification Accuracy, the Confusion Matrix, and the Area Under the ROC Curve (AUC).
Classification accuracy is defined as the ratio of correctly predicted labels to the total number of predictions:
$\text{Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total predictions}},$
with values ranging between 0 and 1. It is most appropriate when classes are balanced and misclassifications are equally important.
The confusion matrix provides a more detailed view of performance, especially in multi-class settings. Figure 10 shows an example of a confusion matrix, taken from Akhmetali et al. (2024), who trained a CNN model to distinguish between 5 classes of variable stars. Each row corresponds to the true class, and each column to the predicted class. Ideally, the matrix is diagonal, indicating perfect classification. Normalizing the matrix removes dependence on class size, allowing for easier comparison across categories.
Figure 10: Example of a confusion matrix taken from Akhmetali et al. (2024), who trained a CNN model to distinguish between 5 classes of variable stars. The confusion matrix displays the percentage distribution of true class labels against the predicted class labels. In the best-case scenario, the confusion matrix is purely diagonal, with all off-diagonal elements equal to zero, indicating that all instances were correctly classified.
The Receiver Operating Characteristic (ROC) curve is a standard diagnostic tool for binary classification. It plots the True Positive Rate (TPR) against the False Positive Rate (FPR):
$\text{TPR}=\frac{\text{True Positives}}{\text{Total Actual Positives}},$
$\text{FPR}=\frac{\text{False Positives}}{\text{Total Actual Negatives}}.$
By varying the classification threshold, the ROC curve visualizes model performance across different operating points. The Area Under the Curve (AUC) provides a summary measure of overall performance, with higher AUC values indicating better discrimination. ROC curves are valuable for comparing models and for guiding the selection of classification thresholds that balance sensitivity and specificity, as illustrated in Figure 11.
In multiclass settings, ROC analysis requires adaptation because TPR and FPR are inherently defined for binary outcomes. One approach is the one-vs-rest (OvR) strategy, in which each class is treated as the positive class while all other classes are treated as negative, producing a separate ROC curve for each class. Alternatively, macro-averaging computes the AUC for each class and averages them equally, giving each class the same weight, while micro-averaging aggregates true positives, false positives, and false negatives across all classes before calculating TPR and FPR, effectively weighting classes by prevalence. Interpreting ROC results in multiclass problems can be challenging, particularly when dealing with imbalanced datasets or more than three classes, and visualizing multiple ROC curves simultaneously may be difficult.
As an alternative to ROC, Precision-Recall (PR) curves are often more informative for imbalanced datasets because they emphasize performance on the positive class. PR curves can also guide threshold selection by highlighting points that optimize the trade-off between precision and recall.
By analyzing ROC or PR curves, it is possible to choose classification thresholds that optimize specific metrics such as the F1 score, balanced accuracy, or sensitivity, rather than relying on a default threshold of 0.5. This is especially important in astronomy, where rare events, such as supernovae or transient detections, require prioritizing sensitivity to avoid missing important signals.
Figure 11: An illustration of a ROC curve taken from Baron (2019), which plots the TPR against the FPR. Ideally, a model should achieve a TPR of 1 with a FPR of 0. The black dashed line represents the performance of a random classifier. By adjusting the model’s hyperparameters, different points on the ROC curve are generated using validation set results. The pink curve corresponds to model 1, with points A, B, and C indicating specific hyperparameter settings. The purple curve shows the performance of model 2. The area under the ROC curve (AUC) provides a measure for model comparison - model 1 performs better in this case.
4.5.2 Validation Methods
To reliably assess the generalization ability of ML models, it is crucial to evaluate them on data that was not used during training. This is achieved through various validation methods, which help prevent overfitting and provide unbiased estimates of model performance.
Holdout Validation (HV) Kim (2009) is the most common approach for evaluating ML models. In this approach, the dataset is split into training, validation, and test sets. A typical split uses 70% for training, 15% for validation, and 15% for testing, though these ratios may vary depending on dataset size and variability. For small datasets, a larger training proportion is preferred to ensure sufficient learning. As datasets grow, smaller validation sets can still capture data diversity effectively. The training set is used to fit model parameters, while the validation set is used for hyperparameter tuning—a process that, although computationally intensive, can be automated and parallelized. The final model is evaluated on the unseen test set. HV is used in various astronomical tasks, such as classification of spectral data Solorio-Ram´ırez et al. (2023), estimation of galaxy redshifts Ball et al. (2008), classification of sterllar variability Audenaert et al. (2021), and predicting stellar rotations Gomes et al. (2024).
Cross-validation (CV) Stone (1974) is a resampling technique used to obtain an unbiased estimate of a model’s performance. Compared to simple holdout validation, it generally offers a more reliable assessment of a model’s generalization error, particularly when working with small datasets. Several cross-validation strategies commonly used for model evaluation are outlined below. CV is used in astronomy for estimation of photometric redshifts of quasars Ball et al. (2007), Star–Galaxy–QSO classification Bai et al. (2018), calibration of ages for the APOGEE DR17 catalogue Boulet (2024), and classification of variabile sources Richards et al. (2012b).
K-Fold Cross-Validation (KFCV) Bengio & Grandvalet (2004) partitions the dataset randomly into $k$ disjoint subsets, or folds. In each iteration, one fold is held out as the validation set, while the remaining k1k-1 folds are used for training. This process repeats $k$ times, ensuring each fold serves exactly once as the validation set. The average of the performance metrics across all iterations provides a robust estimate of the model’s validation error. KFCV is used in various works, such as classification of variable stars Akhmetali et al. (2024), star–galaxy separation Machado et al. (2016b), protostellar classification Miettinen (2018), and identification and fitting eclipse maps of exoplanets Hammond et al. (2024).
Stratified K-Fold Cross-Validation (SKFCV) Prusty et al. (2022) is particularly useful in classification tasks with imbalanced class distributions. Unlike standard KFCV, stratified KFCV preserves the original class proportions within each fold. This ensures that minority classes are adequately represented, yielding more stable and representative performance estimates. SKFCV is used for prediction of habitability of exoplanets Raminaidu et al. (2023), classification of massive stars in nearby galaxies Maravelias et al. (2022), spectral classification of OB stars Kyritsis et al. (2022), and detection of quasi-periodic oscillations in X-ray binaries Kiker et al. (2023).
Leave-One-Out Cross-Validation (LOOCV) Wong (2015) is a special case of KFCV where $k=n$, the total number of data points. Each iteration uses a single sample for validation and the remaining n1n-1 samples for training. LOOCV provides nearly unbiased performance estimates, making it well-suited for very small or imbalanced datasets. However, it is computationally intensive and often impractical for large datasets or models with long training times. SKFCV is for exoplanet atmospheric analysis Welbanks et al. (2023), selection of optically variable active galactic nuclei De Cicco et al. (2025), analysis of stellar LCs Pan et al. (2024), and exoplanet detection Pimentel et al. (2024).
Nested Cross-Validation (NCV) Varma & Simon (2006) addresses the challenge of reliable hyperparameter tuning and unbiased performance estimation. The data is split into outer and inner loops of cross-validation. The inner loop is used for model selection (e.g., tuning hyperparameters), while the outer loop assesses the generalization error. This method helps avoid optimistic bias that can occur when hyperparameters are tuned and evaluated on the same data. NCV is used for star/galaxy classification Chao et al. (2020), cosmological inference from non-linear scales Yuan et al. (2024), classification of variable stars Long et al. (2012a), and prediction of galaxy metallicity from three-colour images Wu & Boada (2019).
Validation methods not only influence model selection but also help assess uncertainty, bias, and variance. For instance, large variations in performance across folds may indicate that the model is sensitive to the training data and prone to overfitting. Similarly, if the model consistently underperforms across validation sets, it may be underfitting or affected by biased assumptions.
Despite the common use of performance metrics and validation schemes, explicit statistical comparisons between models are still uncommon in astronomy. This is problematic because reported differences in metrics (e.g., accuracy or AUC) may not be statistically significant, especially with noisy or imbalanced data. Statistical significance tests help assess whether performance improvements are robust rather than due to random fluctuations Demšar (2006). For instance, McNemar’s test McNemar (1947) evaluates paired classification outcomes, while paired $t$-tests or Wilcoxon signed-rank tests Higgins (2004) can compare cross-validation scores. Bootstrap and permutation methods can also provide confidence intervals for performance differences Good (2005). Incorporating such tests offers a more rigorous foundation for model selection and helps prevent overstatement of results.
5 Applications of Machine Learning in Light Curve Analysis
The increasing data volume from large-scale astronomical surveys such as Kepler, TESS, and the upcoming LSST presents both unprecedented opportunities and significant challenges in processing and interpretation. Traditional methods struggle with the scale, complexity, and noise in LC data, making automation essential. ML has emerged as a transformative approach, enabling efficient classification, detection, and characterization of astronomical objects with remarkable accuracy.
ML techniques are particularly effective at handling the high-dimensionality and noise inherent in LC data. By leveraging SL, UL, and SSL algorithms, ML facilitates pattern recognition, object classification, and anomaly detection with greater efficiency and precision than traditional methods. This section explores three pivotal applications of ML in LC analysis: (1) exoplanet detection, (2) variable star analysis, and (3) supernova classification.
5.1 Transiting Exoplanet Detection
Transit photometry has become a fundamental technique for exoplanet discovery, detecting exoplanets by identifying periodic dips in a star’s brightness and serving as a cornerstone of modern planet-hunting surveys. Figure 12 illustrates a typical LC with a transit event, showcasing the subtle flux dip that ML algorithms are trained to recognize.
Figure 12: Example of transiting exoplanets. The figure is sourced from Shallue & Vanderburg (2018).
Space-based missions including Kepler (2009-2018) and TESS (2018-present) have revolutionized this field by providing high-precision photometric data across large sky areas. The Kepler mission identified thousands of planetary candidates through continuous monitoring of a single field, while TESS has expanded this catalog using its all-sky survey strategy focused on brighter stars. The analysis of these datasets presents significant computational challenges due to their volume, noise characteristics, and the presence of astrophysical false positives, necessitating advanced analytical approaches. Table LABEL:table2 presents a summary of ML methods applied to different exoplanet detection datasets.
The Kepler mission data have served as a testbed for developing ML techniques in transit detection. Initial work established the effectiveness of CNNs through the AstroNet architecture Shallue & Vanderburg (2018), which achieved classification performance comparable to human experts. Subsequent developments introduced modifications such as ExoNet Ansdell et al. (2018), incorporating additional diagnostic information to reduce false positives, and AstroNet-K2 Dattilo et al. (2019), adapted for the modified observing strategy of the K2 mission. Alternative approaches including 2D-CNN architectures Chintarungruangchai & Jiang (2019) and ensemble methods Priyadarshini & Puri (2021) demonstrated improved sensitivity to low signal-to-noise transits.
Analysis of TESS data has built upon these foundations while addressing the mission’s distinct characteristics. Modified versions of the original AstroNet framework have been applied to TESS observations Yu et al. (2019), with subsequent refinements such as Astronet-Triage-v2 Tey et al. (2023) improving classification accuracy. The Nigraha pipeline Rao et al. (2021) represents a comprehensive implementation integrating multiple analysis stages, while systems like SHERLOCK Dévora-Pajares et al. (2024) provide end-to-end processing capabilities. The ExoMiner Valizadegan et al. (2022) and its enhanced version ExoMiner++  Valizadegan et al. (2025) have demonstrated the potential of DL to replicate and augment expert vetting processes.
ML applications have been successfully adapted to various astronomical surveys and simulated datasets beyond the primary space-based missions. In simulated LC analysis, CNNs were applied to detect transits of habitable planets in high-cadence data Zucker & Giryes (2018), while alternative approaches included 1D CNNs for processing non-phase-folded LCs Iglesias Álvarez et al. (2023). GPU-accelerated phase-folding algorithms were developed specifically for detecting ultrashort-period exoplanets Wang et al. (2024).
For ground-based surveys, different ML approaches were implemented: RF and SOM techniques were combined in the NGTS survey for candidate vetting Armstrong et al. (2018), and CNNs were employed for automated candidate screening Chaushev et al. (2019). The QES project utilized DBSCAN-based algorithms for effective noise rejection in transit data Mislis et al. (2018). Similarly, the WASP survey integrated RF and CNN methods for comprehensive transit signal analysis Schanche et al. (2019).
CNNs were also implemented for the BRITE mission’s photometric data analysis Yeh & Jiang (2020). For infrared observations, LSTM networks were applied to Spitzer data for improved detrending of LCs Morvan et al. (2020). In Gaia photometry, XGBoost-assisted methods were developed for transit searches, leading to confirmed exoplanet discoveries Panahi et al. (2022).
These diverse applications demonstrate the versatility of ML techniques across various observational platforms, data types, and specific scientific requirements. The systematic development of analysis methods, reflects the ongoing evolution of techniques to address the challenges posed by current and future transit surveys. These methodological advances continue to enhance the detection and characterization of exoplanetary systems across diverse observational datasets.
Table 2: Summary of ML methods applied to exoplanet detection datasets.
Data type Method Description
Kepler CNN Pearson et al. (2018) proposed a CNN-based method for exoplanet detection, outperforming least-squares techniques without requiring model fitting.
Shallue & Vanderburg (2018) developed AstroNet, a deep CNN designed for exoplanet classification.
Ansdell et al. (2018) proposed ExoNet, extending AstroNet by incorporating domain knowledge, centroid time-series data, and stellar parameters.
Dattilo et al. (2019) proposed AstroNet-K2, an extension of AstroNet adapted for Kepler’s K2 data.
Chintarungruangchai & Jiang (2019) proposed a 2D-CNN model with phase-folding for transit detection, demonstrating improved accuracy at low S/N.
Priyadarshini & Puri (2021) proposed an Ensemble-CNN model for exoplanet detection, comparing its performance with various ML algorithms.
Bugueño et al. (2021) proposed a CNN-based exoplanet detection method using MTF to transform unevenly sampled LCs into fixed-size images.
Cuéllar et al. (2022) proposed a CNN-based transit detection model trained on mixed real and synthetic data.
RF Jenkins et al. (2012) proposed a RF-based approach to automate transit signal classification, generating a preliminary list of planetary candidates.
McCauliff et al. (2015) expanded RF-based exoplanet classification by transforming transit-like detections into numerical attributes.
Sturrock et al. (2019) developed an RF-based exoplanet classification model and deployed it as a publicly accessible API in the cloud.
Caceres et al. (2019) developed the ARPS method combining ARIMA modeling, transit comb filtering, and RF classification to identify exoplanet candidates.
Jin et al. (2022) optimized SL with feature selection and tuning, with RF performing best, and used clustering to identify potentially habitable exoplanets.
Hesar & Foing (2024) evaluated six classification algorithms for exoplanet detection, identifying RF and SVM as the top performers based on accuracy and F1 score.
KNN Thompson et al. (2015) proposed a ML-based metric using dimensionality reduction and KNN to identify transit-shaped signals.
Bahel & Gaikwad (2022) explored exoplanet detection using ML classification, applying KNN on SMOTE-balanced data.
Ensemble Hesar et al. (2024) applied ML models to estimate stellar rotation periods, demonstrating that Voting Ensemble improves accuracy over traditional approaches.
Luz et al. (2024) evaluated five Ensemble ML algorithms for exoplanet classification.
ANN Kipping & Lam (2016) developed an ANN-based model to predict short-period transits likely to have additional planets.
SOM Armstrong et al. (2016) developed a SOM-based method for fast exoplanet candidate classification.
GPC Armstrong et al. (2021) proposed a GPC-based probabilistic planet validation method as an alternative to VESPA.
LightGBM Malik et al. (2022) proposed a ML approach using TSFresh-extracted features and a gradient boosting classifier for transit detection.
GAN Suresh et al. (2024) explored GAN-based data augmentation for exoplanet detection, showing comparable accuracy with synthetic data and improved performance.
TESS CNN Yu et al. (2019) modified AstroNet for automated triage and vetting of TESS candidates.
Osborn et al. (2020) adapted ExoNet for TESS data, training on simulated LCs.
Rao et al. (2021) developed Nigraha, built upon AstroNet, a pipeline combining transit detection, supervised ML, and vetting.
Olmschenk et al. (2021) developed a CNN for efficient exoplanet transit detection, and identified 181 new exoplanet candidates.
Tey et al. (2023) developed Astronet-Triage-v2, built upon AstroNet, an improved neural network for exoplanet candidate triage.
Fiscale et al. (2023) demonstrated that combining transfer learning with regularization techniques significantly enhances CNN performance.
Liao et al. (2024) proposed a wavelet-transform-based LC representation and an improved Inception-v3 CNN.
Dévora-Pajares et al. (2024) proposed SHERLOCK, an end-to-end pipeline that enables efficient exoplanet searches.
DNN Valizadegan et al. (2022) proposed ExoMiner, a DL classifier that mimics expert vetting for transit signals.
Valizadegan et al. (2025) introduced ExoMiner++, an enhanced version of ExoMiner, improving transit signal classification by integrating transfer learning.
Transformer Salinas et al. (2025) proposed a Transformer-based NN for exoplanet detection, identifying transit signals without phase folding or periodicity assumptions.
Simulated data CNN Zucker & Giryes (2018) proposed a CNN-based approach to detect transits of habitable planets in simulated high-cadence LCs.
Iglesias Álvarez et al. (2023) developed a 1D CNN for detecting transits in non-phase-folded LCs.
Wang et al. (2024) proposed GPFC, a GPU-accelerated phase-folding algorithm for ultrashort-period exoplanet detection.
NGTS RF and SOM Armstrong et al. (2018) developed autovet, a ML pipeline, combining RFs and SOMs to rank planetary candidates with high accuracy.
CNN Chaushev et al. (2019) applied a CNN for automated vetting of exoplanet candidates, reducing manual effort.
QES DBSCAN Mislis et al. (2018) developed TSARDI, a DBSCAN-based UL algorithm for noise rejection in transit surveys.
WASP RF and CNN Schanche et al. (2019) developed a ML pipeline combining RFs and CNNs for automated vetting of transit signals.
BRITE CNN Yeh & Jiang (2020) applied CNNs to BRITE LCs for exoplanet transit detection.
Spitzer Space Telescope LSTM Morvan et al. (2020) proposed TLCD-LSTM, a probabilistic LSTM-based detrending method for transit LCs.
Gaia XGBoost Panahi et al. (2022) developed a ML-assisted transit search in Gaia photometry, leading to the first exoplanet detections by Gaia, confirmed as hot Jupiters via radial velocity measurements.
5.2 Variable Star Analysis
Stellar variability is a fundamental characteristic observed in numerous stars across the optical band, manifesting as periodic, semi-regular, or completely irregular brightness fluctuations. Variable stars are broadly categorized into intrinsic and extrinsic variables based on the underlying mechanisms driving their luminosity variations.
Table 3: Variable star types and their corresponding abbreviations Yu et al. (2021).
Variable star types Abbreviation
Eclipsing binary: Algol type EA
Eclipsing binary: Beta type EB
Eclipsing binary: W Ursae Majoris type EW
Ellipsoidal binaries ELL
Long period variable LPV
Mira MIRA
RV Tauri RV
W Virginis: period <<8 d CWA
W Virginis: period >>8 d CWB
RS Canum Venaticorum RS
BY Draconis BY
Population II Cepheid PTCEPH
Delta Cepheid DCEP
Delta Scuti DSCT
Gamma Doradus GDOR
B emission-line star BE
Gamma Cassiopeiae GCAS
Alpha Cygni ACYG
Beta Cephei BCEP
Alpha-2 Canum Venaticorum ACV
RR Lyrae: RRab type RRAB
RR Lyrae: RRC type RRC
RR Lyrae: RRd type RRD
Slowly pulsating B star SPB
Intrinsic variables experience genuine changes in luminosity due to internal physical processes, such as stellar pulsations, eruptions, or structural expansion and contraction. This category primarily includes pulsating and eruptive variables. Pulsating variables undergo periodic expansions and contractions in their outer layers, leading to observable brightness oscillations. Notable examples include Cepheids, RR Lyrae, and Mira variables, each exhibiting distinct pulsation periods and amplitude variations. Eruptive variables, such as cataclysmic variables and nova-like stars, exhibit sudden and often dramatic changes in brightness, typically due to stellar outbursts or accretion-related instabilities.
Figure 13: Variable star classification hierarchy Yu et al. (2021).
Extrinsic variables, in contrast, exhibit brightness fluctuations due to external factors, such as eclipses or rotational modulations. This category encompasses eclipsing binaries and rotating variables. Eclipsing binary systems consist of two or more gravitationally bound stars orbiting a common center of mass, where periodic eclipses result in characteristic minima in their LCs. The primary and secondary minima in their LCs provide insights into stellar radii, temperatures, and orbital inclinations. Rotating variables exhibit modest brightness variations arising from stellar surface features, such as starspots or ellipsoidal distortions, modulated by the star’s rotation.
Figure 13 presents a schematic classification of variable stars, while Table 3 lists their primary types and corresponding abbreviations. The analysis of variable star LCs plays a crucial role in distinguishing between different classes. Pulsating variables, for instance, exhibit smooth and regular brightness oscillations, whereas eclipsing binaries show well-defined periodic minima. These distinctions enable robust classification and facilitate astrophysical inferences about stellar structure, evolution, and binary interactions.
The systematic classification of variable stars relies heavily on distinguishing subtle features in their LCs, such as periodicity, amplitude, and morphological patterns. Traditional classification frameworks often depend on manually engineered features (e.g., periodograms, Fourier coefficients, or phased-folded curve statistics), which may fail to capture nuanced or non-linear relationships in large datasets. ML methods overcome these limitations by automating feature extraction and enabling robust classification across diverse variable star populations. Table LABEL:table4 provides a comprehensive summary of ML techniques applied to variable star analysis, reflecting the evolution of methodologies from early Bayesian approaches to modern DL architectures.
Early ML implementations focused on probabilistic methods and ensemble techniques. Bayesian Networks (BNs) and Gaussian Mixture Models (GMMs) were employed for CoRoT and Hipparcos datasets to probabilistically associate LC features with physical classes Debosscher et al. (2007); Sarro et al. (2006). The introduction of RFs marked a significant advancement, enabling feature importance analysis and improved handling of imbalanced datasets Richards et al. (2011a); Dubath et al. (2011). Subsequent hybrid approaches combined RFs with dimensionality reduction techniques like PCA or SOMs to enhance interpretability Armstrong et al. (2015); Rimoldini et al. (2012).
The advent of DL revolutionized variable star classification by leveraging raw or minimally preprocessed LCs. CNNs achieved state-of-the-art performance on surveys like CRTS and Kepler, identifying hierarchical patterns directly from flux measurements Mahabal et al. (2017); Akhmetali et al. (2024). RNNs and LSTM networks proved particularly effective for capturing temporal dependencies in irregularly sampled data from ASAS-SN and Gaia Naul et al. (2018); Merino et al. (2024). Transformer architectures, recently applied to Kepler and ZTF data, demonstrated superior performance in modeling long-range dependencies without phase-folding assumptions Pan et al. (2024); Cádiz-Leyton et al. (2024).
Table 4: Summary of ML methods applied to variable star analysis.
Data type Method Description
ASAS Bayesian Classifier Eyer & Blake (2005) developed a Fourier-based Bayesian classifier for ASAS variables.
Hipparcos Bayesian ensemble Sarro et al. (2006) developed a Bayesian ensemble of neural networks for automatic classification of eclipsing binary LCs.
OGLE BN and SVM Debosscher et al. (2007) presented BN and SVM in the application of the methodology to variable stars.
OGLE BN and SVM Sarro et al. (2009) developed and tested a Fourier-based Bayesian classifier
CoRoT BN and GMM Debosscher et al. (2009) developed a fast pipeline for classifying CoRoT LCs and discovering new stellar variability types.
OGLE and Hipparcos RF Richards et al. (2011a) developed a ML methodology using RF.
Hipparcos RF and BN Dubath et al. (2011) evaluated automated classification of Hipparcos periodic stars using RFs.
Hipparcos RF and BN Rimoldini et al. (2012) applied RFs to classify periodic, non-periodic, and irregular Hipparcos variables.
Kepler RF Long et al. (2012b) introduced noisification to reduce survey-dependent feature mismatch.
Hipparcos and OGLE Active Learning (AL) Richards et al. (2011b) used active learning to reduce sample bias in variable star classification
PTF RF Bloom et al. (2012) developed a ML framework for PTF to automate discovery and classification of transients and variables.
LINEAR GMM and SVM Sesar et al. (2013) proposed method for identifying visually confirmed variable stars within the LINEAR survey.
LINEAR GMM and SVM Palaversa et al. (2013) identified $\sim$7000 faint periodic stars using LINEAR and SDSS data.
MACHO BN and RF Nun et al. (2014) proposed methodology for anomaly detection in MACHO data.
WISE RF and AL Masci et al. (2014) proposed methodology for classifying periodic variable stars.
Catalogue of Eclipsing Variables Membership probability Avvakumova & Malkov (2014) developed a procedure to classify eclipsing binaries based on LC parameters.
OGLE and ASAS KNN, SVM and RF Kügler et al. (2015) proposed a density model for classifying irregular time-series data.
Kepler SOM and RF Armstrong et al. (2015) developed a novel method for classifying variable stars by combining SOM and RF.
Kepler RF and BN Bass & Borne (2016) proposed an ensemble approach for variable star classification in the Kepler field.
MACHO, LINEAR and ASAS RF Kim & Bailer-Jones (2016) developed a general-purpose ML package for classifying periodic variable stars.
MACHO and OGLE SVM Mackenzie et al. (2016) extracted subsequences of LCs and clustered them to identify common local patterns.
UCR and LINEAR KNN, RF, RBF-NN Johnston & Peter (2017) developed a novel time-domain feature extraction method called Slotted Symbolic Markov Modeling (SSMM).
ASAS, Hipparcos and OGLE KNN, SVM, and RF Johnston & Oluseyi (2017) proposed a method for classifying variable stars using supervised pattern recognition.
CRTS CNN Mahabal et al. (2017) developed a DL approach for classifying LCs by transforming sparse, irregular time-series data into 2D representation.
ASAS, LINEAR, MACHO RNN Naul et al. (2018) developed an unsupervised autoencoding RNN that effectively handles irregularly sampled, noisy LCs.
Kepler LSTM and RNN Hinners et al. (2018) presented methods for representation learning and feature engineering aimed at predicting and classifying properties.
OGLE LR, SVM, KNN, RF, and SGB Pashchenko et al. (2018) proposed a ML approach for variability detection.
OGLE, MACHO and Kepler Fast Similarity Function Valenzuela & Pichara (2018) presented a novel data structure called the Variability Tree.
Gaia and ASAS RF Jayasinghe et al. (2019) utilized the RF classifier along with a series of classification corrections.
OGLE, VISTA and CoRoT CNN Aguirre et al. (2019) developed a scalable CNN architecture for survey-independent LC classification.
ASAS-SN RNN Tsang & Schultz (2019) developed classifier combining an RNN AE with a GMM.
ASAS PCA and RF McWhirter et al. (2019) focused on processing time-series data with uneven cadence by leveraging representation learning to extract useful features.
CRTS RF Hosenie et al. (2019) developed an optimized ML framework for variable star classification.
Light curves in Galactic Plane $k$-medoids method Modak et al. (2020) proposed a $k$-medoids clustering approach to objectively classify galactic variable stars.
CoRoT, OGLE and MACHO Streaming Probabilistic Model Zorich et al. (2020) proposed a streaming probabilistic classification model that uses a novel set of features.
OGLE, Gaia and WISE RNN Becker et al. (2020) developed an end-to-end DL approach using RNNs for efficient variable star classification.
CRTS RF and XGBoost Hosenie et al. (2020) proposed a hybrid approach combining HC with data augmentation techniques.
MACHO RNN Jamal & Bloom (2020) conducted systematic comparison of neural network architectures for time-series classification.
UCR Starlight and LINEAR Multi-View Metric Learning Johnston et al. (2020) introduced a Multi-View Metric Learning framework that leverages multiple data representations.
OGLE iTCN and iResNet Zhang & Bloom (2021) developed Cyclic-Permutation Invariant Neural Networks that achieve state-of-the-art accuracy.
OGLE CNN and LSTM Bassi et al. (2021) proposed 1D CNN-LSTM hybrid network for direct variable star classification using raw time-series data.
Simulated data LSTM Čokina et al. (2021) developed a DL method for automated classification of eclipsing binaries.
ZTF BRF Sánchez-Sáez et al. (2021) introduced ALeRCE’s first LC classifier, a two-level balanced RF system processing ZTF alerts.
Kepler GMM Barbara et al. (2022) developed an interpretable classification system for Kepler LCs.
OGLE Multiple-Input Neural Network Szklenár et al. (2022) developed a Multiple-Input Neural Network combining CNNs and MLPs.
OGLE, CSS, Gaia UMAP and HDBSCAN Pantoja et al. (2022) developed semi-supervised and clustering-based approaches for variable star classification.
VVV RF and XGBoost Molnar et al. (2022) developed VIVACE, an automated two-stage classification pipeline.
ZTF CVAE Chan et al. (2022) proposed an unsupervised DL approach using variational AE and isolation forests.
Kepler ResNet and LSTM Yan et al. (2023) developed RLNet, a hybrid ResNet-LSTM neural network.
TESS SVM Elizabethson et al. (2023) developed a ML framework classifying T Tauri stars into 11 morphological classes.
Gaia LSTM and GRU Merino et al. (2024) proposed a self-supervised learning approach using RNNs.
LAMOST LightGBM and XGBoost Qiao et al. (2024) developed a LightGBM/XGBoost-based classification system for LAMOST DR9 data.
OGLE CNN Monsalves et al. (2024) developed an efficient CNN-based classification system using 2D histogram representations of OGLE LCs.
TESS CNN Olmschenk et al. (2024) developed a rapid CNN classifier for TESS 30-minute cadence data.
TMTS XGBoost and RF Guo et al. (2024) developed a classification system for TMTS variables using XGBoost and RF.
ZTF Distance Metric Classifier Chaini et al. (2024) developed DistClassiPy, an interpretable distance-metric classifier for variable stars.
MACHO, OGLE and ATLAS Transformer Cádiz-Leyton et al. (2024) proposed HA-MC Dropout, a novel transformer-based method combining hierarchical attention and Monte Carlo dropout.
Kepler Transformer Pan et al. (2024) developed Astroconformer, a Transformer- based model that demonstrates superior performance.
OGLE CNN Akhmetali et al. (2024) developed a CNN-based approach for automated variable star classification.
5.3 Supernova classification
Supernovae (SNe) are among the most energetic transient phenomena in the universe. They play a critical role in stellar evolution, the chemical enrichment of the interstellar medium, and cosmological distance measurements. Their classification traditionally relies on spectroscopic and photometric observations. While spectroscopic classification remains the most accurate, it is expensive and has high requirements for telescopes and observation time, and thus cannot be applied to all observed transients.
Photometric classification, although less precise, offers higher observational efficiency and has gained importance with the advent of wide-field surveys. Early photometric methods used template fitting and parametric modeling of LCs, leveraging features such as peak brightness, decline rate, and color evolution. However, these methods typically require complete LCs with full phase coverage, limiting their application to real-time or sparsely sampled data.
Recent advances in ML have significantly enhanced SN classification capabilities, especially under constraints such as low signal-to-noise ratios and incomplete data. ML-based approaches can classify a wide range of transient types and support near real-time classification. This is critical for follow-up prioritization and maximizing the scientific return from transient surveys. Table 5 provides SN classes and their physical origins, while Table LABEL:table6 summarizes various ML methods applied to SN classification.
Table 5: Major supernova classes and their physical origin Gal-Yam (2016).
Supernova types Physical origin
SN Ia White Dwarf
SN Ib Massive star
SN Ic Massive star
SN Ic-BL Massive star
SN II Massive star
SN IIb Massive star
SN IIn Massive star
SLSN Massive star
Early ML applications focused on engineered features derived from parametric LC fits or domain-specific metrics. The Supernova Photometric Classification Challenge (SNPCC) Kessler et al. (2010) was created as a standardized dataset to evaluate ML-based photometric classification methods for SN. Many subsequent studies tested their algorithms on the SNPCC dataset Newling et al. (2011); Richards et al. (2012a); Karpenka et al. (2013); Gupta et al. (2016); Lochner et al. (2016); Charnock & Moss (2017); Ishida et al. (2019); Pasquet et al. (2019); Santos et al. (2020); de Oliveira et al. (2023).
DL revolutionized SN classification by allowing models to process raw flux measurements or minimally preprocessed LCs. Initial ML applications employed a variety of feature‐based techniques: Kernel Density Estimation (KDE) and Boosting methods Newling et al. (2011), Non‐linear Dimensionality Reduction with RFs Richards et al. (2012a), ANNs Karpenka et al. (2013), Domain Adaptation with AL Gupta et al. (2016), and Naive Bayes, KNN, SVM, ANN, and BDTs Lochner et al. (2016).
RNN and CNN architectures then advanced photometric classification by ingesting raw time series directly. Deep RNNs (including LSTM variants) demonstrated strong performance on SNPCC and SALT2‐fitted LCs Charnock & Moss (2017); Möller & de Boissière (2020), while CNNs learned hierarchical features from 2D representations of LCs Brunel et al. (2019); Qu et al. (2021). Hybrid models, such as SuperNNova Möller & de Boissière (2020), SuperRAENN Villar et al. (2020), and PELICAN framework Pasquet et al. (2019) leveraged SSL and AEs to boost purity and completeness.
Recent classification pipelines increasingly incorporate generative models, Gaussian Process (GP) augmentation, and real-time alert systems Martínez-Palomera et al. (2022); Boone (2021); Smith & Geach (2023). For example, the ParSNIP framework Boone (2021) utilizes Variational Autoencoders (VAEs) to perform classification. Similarly, Avocado Boone (2019) applies LightGBM combined with GP augmentation to classify transients photometrically. Real-time alert brokers like Fink Leoni et al. (2022) streamline the early identification of SNe from surveys like ZTF using AL strategies. CNN-based models such as SCONE Qu et al. (2021) apply 2D GP regression to multi-band LCs, while Photo-zzSNthesis Qu & Sako (2023) uses CNNs to generate full redshift probability distributions. Temporal convolutional networks (TCNs) paired with LightGBM, as in the TLW model Li et al. (2024), further demonstrate the effectiveness of hybrid architectures for robust, survey-independent transient classification.
Table 6: Summary of ML methods applied to supernova classification.
Data type Method Description
SNPCC KDE and Boosting Newling et al. (2011) proposed two classification methods for the application of SNPCC data.
SNPCC Non-linear Dimension Reduction and RF Richards et al. (2012a) proposed the non-linear dimension reduction technique to detect structure in a data base of SNe LCs.
SNPCC ANN Karpenka et al. (2013) presented a method for automated photometric classification of SNe.
SNPCC Domain Adaptation and AL Gupta et al. (2016) presented an adaptive mechanism that generates a predictive model to identify SNe Ia.
SNPCC Naive Bayesian, KNN, SVM, ANN and BDT Lochner et al. (2016) developed a multi-faceted classification pipeline.
SNLS XGBoost Möller et al. (2016) presented a method to photometrically classify SNe Ia.
SNPCC Deep RNN Charnock & Moss (2017) presented deep RNN for performing photometric classification of SNe.
SNPCC CNN Brunel et al. (2019) presented CNN for SNe Ia classification.
SNPCC AL Ishida et al. (2019) developed a framework for spectroscopic follow-up design for optimizing supernova photometric classification.
PS1-MDS SVM, RF and MLP Villar et al. (2019) developed 24 classification pipelines with different feature extraction and data augmentation methods.
PLAsTiCC DNN Muthukrishna et al. (2019) developed RAPID, a novel time-series classication tool for identifying explosive transients.
PLAsTiCC LightGBM Boone (2019) developed Avocado, a software package for classification of transients with GP augmentation.
SNPCC CNN and AE Pasquet et al. (2019) developed PELICAN, an algorithm for the characterization and the classification of SNe LCs.
SALT2 fitted RNN Möller & de Boissière (2020) developed SuperNNova, a framework for photometric classification of SNe.
PS1-MDS RF and RAENN Villar et al. (2020) developed SuperRAENN, a semi-supervised SN photometric classification pipeline.
OSC and ZTF RF Gomez et al. (2020) developed a classification algorithm targeted at rapid identification of a pure sample of SLSN-I.
SNPCC TPOT, XGBoost, AdaBoost, GBoost, EXT, RF Santos et al. (2020) analyzed the performance of boosting and averaging methods for classification of SNe.
SALT2 DNN Takahashi et al. (2020) developed a classification algorithm to classify LCs observed by Subaru/HSC.
PS1-MDS RF Hosseinzadeh et al. (2020) developed Superphot, an open-source classification algorithm for photometric classification of SNe.
PS1-MDS and PLAsTiCC VAE Boone (2021) developed ParSNIP, a hybrid model to produce empirical generative models of transients from data sets of unlabeled LCs.
PLAsTiCC CNN Qu et al. (2021) developed SCONE, a CNN-based classification method using 2D GP regression.
PLAsTiCC CNN Qu & Sako (2022) presented classification results on early SNe LCs from SCONE.
ZTF AL Leoni et al. (2022) developed Fink, a broker for early SNe classification.
Open Supernova Catalog GP Stevance & Lee (2023) explored the application of GP to SNe LCs.
SNPCC XGBoost de Oliveira et al. (2023) developed a linear regression algorithm optimized through automated machine learning (AutoML) frameworks.
PLAsTiCC and ZTF MLP, NF and Bayesian Neural Network Demianenko et al. (2023) examined several ML-based LC approximation methods.
PLAsTiCC and Simulated SDSS-II SN data CNN Qu & Sako (2023) developed Photo-zzSNthesis, a CNN-based method for predicting full redshift probability distributions from multi-band SNe LCs.
ZTF AL Pruzhinskaya et al. (2023) explored the potential of AL techniques in application to detect new SNe candidates.
ZTF LightGBM de Soto et al. (2024) developed Superphot+, a photometric classifier for SNe LCs that does not rely on redshift information.
PLAsTiCC TCN and LightGBM Li et al. (2024) developed TLW, a classification algorithm for transients.
6 Challenges and Open Issues
The application of ML to LC analysis has revolutionized astronomical research, yet significant challenges remain in developing robust, interpretable, and survey-independent classification systems. These challenges are expected to intensify with upcoming data-intensive surveys such as the LSST, which will detect tens of millions of transient events nightly.
A key difficulty lies in the heterogeneity and sparsity of data. Surveys like ASAS-SN and Gaia produce irregularly sampled LCs due to observational constraints such as weather or cadence gaps. This irregularity complicates the use of DL models, which typically assume regularly sampled inputs. Moreover, models trained on one survey (e.g., Kepler) often perform poorly when applied to others (e.g., TESS), due to differences in cadence, noise levels, and photometric filters Valizadegan et al. (2022). These issues are further exacerbated by limited spectroscopic follow-up, especially for rare classes such as SLSNe.
Model interpretability and physical consistency are additional concerns. Many ML models act as black boxes, offering little insight into their predictions. This lack of transparency hinders their adoption in cosmological studies that require rigorous uncertainty quantification and interpretability. Hybrid approaches that incorporate physical priors into data-driven models are a promising direction.
Real-time processing and early classification present another critical challenge. LSST will require sub-minute response times for alert streams, straining even highly optimized models. Furthermore, most classifiers rely on full-phase LCs, whereas early-time classification should operate with incomplete data—often limited to the rising or plateau phases. Building models that can deliver accurate early-stage predictions is an active area of research.
Class imbalance and anomaly detection remain persistent issues. Rare transients such as kilonovae and luminous red novae are difficult to detect using conventional techniques. Oversampling strategies often fall short, prompting interest in generative and data augmentation methods. Moreover, most classifiers assume a closed-set of known classes, yet upcoming surveys will likely discover entirely new types of transients.
Generative models—such as GANs, VAEs, and diffusion models—are emerging as powerful tools for addressing these challenges. They enable synthetic augmentation of rare classes, imputation of missing or irregular data, and cross-survey domain adaptation. While still in early development, generative approaches show strong potential to improve the robustness and generalizability of ML-based LC pipelines.
7 Conclusions
The rapid advancement of observational capabilities in astronomy has led to an exponential increase in the volume of LC data, opening up both exciting opportunities and complex challenges for time-domain astronomy. In this evolving landscape, ML has emerged as a powerful tool, finding applications across a broad spectrum of tasks. This review discusses major photometric surveys that provide the essential LC data, outlines the fundamental principles of ML, and explores ML applications in LC analysis, including exoplanet detection, variable star analysis, and supernova classification, highlighting the increasing sophistication and versatility of these methods.
As astronomical surveys scale up in depth, cadence, and volume, the need for automated, scalable, and interpretable analysis pipelines becomes ever more urgent. ML models, particularly DL architectures, have shown exceptional performance in handling large, noisy, and irregular datasets. Importantly, the choice of ML approach depends on the specific scientific goal, the nature of the dataset, and the balance among performance, interpretability, and computational cost. At the same time, critical challenges such as survey dependence, class imbalance, interpretability, and real-time applicability remain open issues.
Future advancements are likely to involve approaches that go beyond purely data-driven methods by integrating physical models, enhancing generalizability across different surveys, and incorporating robust uncertainty quantification.The convergence of domain expertise and machine intelligence holds the key to not only improving classification accuracy but also enabling new scientific insights.
In conclusion, the fusion of ML with astronomical time-series data is not just a technical advancement—it represents a paradigm shift in how discoveries are made. As datasets continue to expand, so too will the opportunities for ML to illuminate the dynamic universe in ways previously unimaginable.